# Distinguishing Tidal Disruption Events and Changing-look Active Galactic Nuclei via Variation of Mid-infrared Color

### v2

@@ Appendix D The parameters of MIR color variation for the MIR selected samples
SDSS J0103+1401 0.04181 Star-forming $-0.07\pm 0.03$ $1.04\pm 0.02$ 0.00% 0.00% 100.00%
@@
SDSS J0859+0922 0.15188 Star-forming $0.07\pm 0.06$ $0.59\pm 0.08$ 0.00% 32.89% 67.11%
@@
SDSS J1152+4850 0.15102 Composite $-0.08\pm 0.12$ $0.83\pm 0.25$ 0.01% 6.96% 93.03%
SDSS J1153+4037 0.1451 Star-forming $-0.02\pm 0.26$ $0.79\pm 0.11$ 5.57% 27.27% 67.15%
@@
SDSS J1448+1137 0.06657 Star-forming $-0.15\pm 0.43$ $1.15\pm 0.38$ 9.40% 17.88% 72.72%
@@
SDSS J1041+3412 0.14028 $-0.20\pm 0.81$ $0.59\pm 0.30$ 23.82% 12.73% 63.44%
SDSS J1051+2101 0.06593 $-0.42\pm 0.61$ $1.07\pm 0.19$ 9.05% 10.80% 80.14%
SDSS J1105+5941 0.03369 $0.11\pm 0.05$ $0.62\pm 0.11$ 0.00% 54.86% 45.14%
@@
SDSS J1430+2303 0.08105 $0.01\pm 0.12$ $0.68\pm 0.24$ 0.04% 22.56% 77.40%
SDSS J1504+0107 0.12826 $0.14\pm 0.16$ $0.52\pm 0.33$ 5.00% 54.47% 40.53%
SDSS J1511+2214 0.12048 $-0.01\pm 0.27$ $0.78\pm 0.20$ 6.51% 27.49% 66.00%
SDSS J1533+2729 0.07193 $-0.18\pm 0.21$ $0.85\pm 0.18$ 0.32% 8.90% 90.77%
@@ Appendix E Simulations of dust echoes
In this section, we described the simulations we made which are not present in the main text. In subsequent simulations, we will focus on the two cases: a TDE case with TDE form UV LCs and no underlying AGN, and a CLAGN case with LINEAR form UV LCs and underlying AGN with $L_{\rm AGN}$ the same as the peak luminosity $L_{\rm max}$.
E.1 Influence by peak UV luminosity and duration
Figure E1: Simulations exploring the influence by peak UV luminosity $L_{\rm max}$ and duration $\tau$. Panels (a) and (c) show the MIR color curves (solid and dashed for the TDE and CLAGN cases, respectively), and (b) and (d) show the inferred ICs and CVRs using the sampled data (circles and squares for the TDE and CLAGN cases, respectively). Different colors indicate different $L_{\rm max}$ for panels (a) and (b), and indicate different $\tau$ for panels (c) and (d).
We studied the influence of different UV peak luminosities $L_{\rm max}$ in the range of $10^{43}-10^{45}$ erg s-1 (Figure E1(a), (b)). For both the TDE and CLAGN cases, CVR in the rising phases decreases with the increase of $L_{\rm max}$, while IC is little affected.
According to the literature, the peak UV luminosities of MIR bright TDEs are between $10^{43.4}$ and $10^{45.7}$ erg s-1 (Onori et al., 2022; Wang et al., 2024; Yao et al., 2023; Earl et al., 2024; Zhang et al., 2025). CLAGNs’ peak UV luminosities range widely. Assuming that $L_{\rm UV}$ is $\sim$10 times $L_{\rm W2}$ according to typical AGNs’ SEDs, the CLAGN sample in this work have peak UV luminosities in the range of $\sim 10^{43}-10^{45}$ erg s-1. There is no significant overall difference in the peak UV luminosities of the TDE and CLAGN samples. Thus, it can hardly explain the observed difference between TDEs and CLAGNs with the difference in peak UV luminosities. However, the dispersion of peak UV luminosities may explain the internal CVR dispersion among TDEs, ANTs or CLAGNs.
We then studied the influence of different durations $\tau$ in the range of $10^{6}-10^{8}$ s (Figure E1(c), (d)). For the TDE case, CVR in the rising phase decreases with the increase of $\tau$. When $\tau$ increases to $\sim 10^{7.5}-10^{8}$ s ($\sim 1-3$ years), the CVR drops to values similar to those in CLAGNs. Fortunately, the durations of most TDEs and ANTs are not that long, and thus the results obtained by assuming $\tau=10^{7}$ s are not affected. While for the CLAGN form, CVR is little affected by $\tau$.
E.2 Influence by shape of dust structure
Figure E2: Simulations exploring the influence of the shape of dust structure. Panels (a), (c) and (e) show the MIR color curves (solid and dashed for the TDE and CLAGN cases, respectively), and (b), (d) and (f) show the inferred ICs and CVRs using the sampled data (circles and squares for the TDE and CLAGN cases, respectively). Different colors indicate different $r_{\rm in}$ for panels (a) and (b), and different $r_{\rm out}$ for panels (c) and (d), and different geometries for panels (e) and (f).
We first studied the influence of the initial inner radius $r_{\rm in}$ on CVR and IC (Figure E2(a), (b)). We tried three $r_{\rm in}$ values of 0.03, 0.1 and 0.3 pc, which are 0.24, 0.8 and 2.4 times $r_{\rm sub}$ ($=0.13$ pc for $L_{\rm max}=10^{44}$ erg s-1), respectively. The results show that $r_{\rm in}$ has little effect on the MIR color variation when it is much smaller than $r_{\rm sub}$. However, when $r_{\rm in}$ is close to or greater than $r_{\rm sub}$, the increase of $r_{\rm in}$ leads to the increase of IC for both TDE and CLAGN cases, and leads to the decrease of CVR for the TDE case.
We then studied the influence of the outer radius $r_{\rm out}$ in the range of 0.3–3 pc (Figure E2(c), (d)). The $r_{\rm out}$ has little effect on CVR, but significantly affects IC. This is because the IR emission in the rising phase mainly comes from the inner region of the dust structure, and the outer region only affects the optical depth of the emergent IR emission. Therefore, the difference in $r_{\rm out}$ cannot explain the difference in CVRs, but can explain the differences and dispersions of ICs. In our simulations, we fixed grain radius $a_{0}$ at 0.1 $\mu$m and number density $n_{d}$ at $10^{8}$ cm-3. The actual quantity directly related to IC is the radial optical depth of the dust shell, which can be described by the dimensionless parameter $\tau_{\rm dust}=(r_{\rm out}-r_{\rm sub})\pi a_{0}^{2}n_{d}$.
We finally studied the influence of different dust geometries between active (torus) and inactive galaxies (roughly isotropic). We further developed the radiative transfer model so that it can calculate the echo of a torus. The torus is described by a half-opening angle and an inclination of $i$ between the axis of the torus and the line of sight. We set the half-opening angle to be $50^{\circ}$, and considered two inclinations, $30^{\circ}$ for face-on torus and $80^{\circ}$ for edge-on torus.
The results are shown in Figure E2(e), (f). The edge-on torus leads to close IC and CVR compared with spherical shell, while face-on torus dust leads to much smaller IC and slightly larger CVR than spherical shell and edge-on torus. The reason for the much smaller IC for face-on torus is that the initial MIR emission, mainly from the inner region of the dust structure, emerges with a much smaller optical depth. The effect would be more pronounced when the dust shell is thicker. In addition, the reason for the slightly larger CVR may be that the echo produced by the peak luminosity is bluer due to smaller emergent optical depth, and has a larger color difference with the echo produced by the long tail and outer region with redder color.
Our radiative transfer model assumes uniform dust, but it is possible that the actual AGN dusty torus is clumpy (Nenkova et al., 2008). Assuming clumpy dust, it can also be deduced that the IR emission from the inner region of face-on torus has a smaller emergent optical depth than those of spherical shell and edge-on torus. Therefore, the above conclusions about IC hold, regardless whether the dust is uniform or clumpy.
E.3 Prediction on variation of NIR color
@@
We studied the possible difference in the NIR color of TDEs and CLAGNs using simulations. For the TDE case, we set an outburst with TDE form UV LC with $L_{\rm max}=10^{44}$ erg s-1. While for the CLAGN case, we set an outburst with LINEAR form UV LC with $L_{\rm max}=5\times 10^{43}$ erg s-1 and an underlying AGN with $L_{\rm AGN}=5\times 10^{43}$ erg s-1. The settings of other parameters were the same as in subsection 4.1.1. We calculated the NIR LCs in H and K bands and the variation of $H-K$ color. As shown in Figure E3, the NIR color of TDEs turns red rapidly in the rising phase, while that of CLAGNs keeps nearly steady in the rising phase. Our simulations suggest that the NIR color variation has the potential to distinguish TDEs from CLAGNs.

### v3

@@ Appendix D The parameters of MIR color variation for the MIR selected samples
SDSS J0103+1401 0.04181 Star-forming $0.07\pm 0.03$ $1.04\pm 0.02$ 0.00% 0.00% 100.00%
@@
SDSS J0859+0922 0.15188 Star-forming $0.07\pm 0.06$ $59\pm 0.08 0.0$0% 32.89% 67.11%
@@
SDSS J1152+4850 0.15102 Composite $0.08\pm 0.12$ $83\pm 0.25 0.0$1% 6.96% 93.03%
SDSS J1153+4037 0.1451 Star-forming $0.02\pm 0.26$ $0.79\pm 0.11$ 5.57% 27.27% 67.15%
@@
SDSS J1448+1137 0.06657 Star-forming $0.15\pm 0.43$ $1.15\pm 0.38$ 9.40% 17.88% 72.72%
@@
SDSS J1041+3412 0.14028 $0.20\pm 0.81$ $0.59\pm 0.30$ 23.82% 12.73% 63.44%
SDSS J1051+2101 0.06593 $0.42\pm 0.61$ $1.07\pm 0.19$ 9.05% 10.80% 80.14%
SDSS J1105+5941 0.03369 $0.11\pm 0.05$ $62\pm 0.11 0.0$0% 54.86% 45.14%
@@
SDSS J1430+2303 0.08105 $0.01\pm 0.12$ $68\pm 0.24 0.0$4% 22.56% 77.40%
SDSS J1504+0107 0.12826 $0.14\pm 0.16$ $2\pm 0.33 5.005$4.47% 40.53%
SDSS J1511+2214 0.12048 $0.01\pm 0.27$ $0.78\pm 0.20$ 6.51% 27.49% 66.00%
SDSS J1533+2729 0.07193 $0.18\pm 0.21$ $0.85\pm 0.18$ 0.32% 8.90% 90.77%
@@ Appendix E Simulations of dust echoes
E.1 Simulation of echo of torus dust
Figure E1: Illustration of the torus geometry in the simulation.
In this subsection, we introduce how we modified Lu16’s model to calculate the echo of torus dust.
The dust geometry is shown in Figure E1. We required that the position $(R,\Theta,\Phi)$ of the dust satisfy $r_{\rm in}<R<r_{\rm out}$ and $\frac{\pi}{2}-\sigma<\Theta<\frac{\pi}{2}-\sigma$, where $r_{\rm in}$ and $r_{\rm out}$ are the inner and outer radii, and $\sigma$ is the half opening angle of the torus. The torus is viewed with an inclination angle $i$, the angle between the line of sight and the torus axis. In this work, we assumed $\sigma=40^{\circ}$, and hence the corresponding covering factor of the torus is 0.65.
The approaches to calculate the echo of torus dust are largely the same as those used to calculate the echo of spherical shell by Lu16. The main differences lie in the following two points. One is that when calculating the total luminosity via volume integration (equation 16 in Lu16), the lower and upper limits of $\theta$ are $\frac{\pi}{2}-\sigma$ and $\frac{\pi}{2}-\sigma$, respectively, instead of 0 and $\pi$. The other is that when calculating the IR emergent optical depth for a point $P$ at a time $t$, instead of using equations 20 and 21 in Lu16, we used the following path integral:
$\tau_{\rm IR}(P,t)=\int_{L}\pi a^{2}(r,t_{r})n_{d}Q_{\rm IR}\Delta(r,\theta){\rm d}l.$ (E1)
In this equation, the path $L$ follows the line of sight. For an infinitesimal path ${\rm d}l$ at a position of $(r,\theta,\phi)$, $a(r,t_{r})$ is the grain size at the position at a delayed time $t_{r}=t+d/c$, where d is the distance to $P$, $\Delta(r,\theta)$ is a function reflecting whether the position is inside or outside the torus:
$\Delta(r,\theta)=\left\{\begin{array}[]{lr}1,\ \ \ \ {\rm if}\ r_{\rm in}<r<{\rm out},\frac{\pi}{2}-\sigma<\theta<\frac{\pi}{2}-\sigma\\ 0,\ \ \ \ {\rm else}\end{array}\right$. (E2)
E.2 Prediction on variation of NIR color
@@
We studied the possible difference in the NIR color of TDEs and CLAGNs using simulations. For the TDE case, we set a flare with TDE form UV LC with $L_{\rm max}=10^{44}$ erg s-1. While for the CLAGN case, we set an outburst with LINEAR form UV LC with $L_{\rm max}=5\times 10^{43}$ erg s-1 and an underlying AGN with $L_{\rm AGN}=5\times 10^{43}$ erg s-1. The settings of other parameters were the same as in subsection 4.1.1. We calculated the NIR LCs in H and K bands and the variation of HKH-K color. As shown in Figure E2, the NIR color of TDEs turns red rapidly in the rising phase, while that of CLAGNs keeps nearly steady in the rising phase. Our simulations suggest that the NIR color variation has the potential to distinguish TDEs from CLAGNs.